# Shadow in the Cache: Unveiling and Mitigating Privacy Risks of KV-cache in LLM Inference

### v1

@@ V-C2 Evaluation of Security
TABLE II: Efficacy of Defense Mechanisms Against Input Reconstruction Attacks.
@@
This section evaluates the capability of different protection schemes to defend against our three KV-cache-based reconstruction attacks. We repeat the attack experiments on the KV-cache after applying each defense mechanism. As shown in Table II (evaluation on the remaining models can be found in Appendix -E), for the “Plaintext” baseline, all three attacks achieve very high success rates. After applying KV-Cloak, the semantic similarity (BERTScore) of all attack outputs drops to a level consistent with random chance, and the ROUGE-L score falls to nearly 0. These results are statistically indistinguishable from comparing the original input with a random string, demonstrating that semantic reconstruction is entirely disrupted. This proves that KV-Cloak effectively protects the private information within the KV-cache. For the DP baseline, its defensive efficacy is strongly correlated with the privacy budget $\epsilon$. With a weak budget of $\epsilon=10^{8}$, the accuracy of the inversion and injection attacks is reduced, but the collision attack can still recover some useful information.
We also compare the distributions of $dist_{target}$ and $dist_{other}$ for the collision attack on the last layer of LLaMA-3.2-1B’s KV-cache under different protections. As shown in Figure 9, the two distributions remain effectively distinguishable under DP protection. When using an enhanced threshold derived from prior knowledge, the per-token attack success rate is as high as 94.84% for ($10^{7},10^{-5}$)-DP, and achieving nearly 100% (the same as plaintext) for ($10^{7},10^{-5}$)-DP. In contrast, under KV-Cloak’s protection, the two distributions become completely indistinguishable, resulting in a 0% attack success rate.
@@
Figure 9: Distributions of $dis_{target}$ (orange) and $dis_{other}$ (blue) for a collision attack on the last-layer KV-cache. The input is an excerpt from “The Bitter Lesson” to the LLaMA-3.2-1B model. The distributions are shown under four conditions: (a) Plaintext, (b) KV-Cloak protection, (c) ($10^{7},10^{-5}$)-DP protection, and (d) ($10^{8},10^{-5}$)-DP protection. The green dotted line is the baseline heuristic threshold ($3\sigma_{other}$), and the red dashed line is the enhanced threshold assuming rank $r=8$.
To assess the robustness of our defense, we also tested KV-Cloak’s performance against the enhanced collision attack on multiple LLMs. As illustrated in Table II, our mechanism demonstrates strong defensive capabilities. With KV-Cloak enabled, the reconstruction accuracy of the attack drops to near-zero, and the recovered outputs are qualitatively equivalent to random noise, confirming that KV-Cloak effectively neutralizes this advanced attack vector.
Takeaway 5: KV-Cloak completely thwarts all our proposed attacks, reducing the quality of any reconstructed text to a level statistically indistinguishable from random noise.
V-C3 Inference Accuracy
To evaluate the model fidelity of our KV-Cloak, we simulate an inference service with a prefill-decode architecture. The KV-cache generated during the prefill phase is protected using either DP or KV-Cloak. This protected cache is then passed to the decode node (after de-obfuscation in KV-Cloak’s case) for subsequent token generation. We then evaluate this protected inference service on the MMLU and SQuAD benchmarks to measure the impact of each method on model accuracy.
@@ V-C3 Inference Accuracy
We tested the impact of KV-Cloak on all experimental models. The results in Table III show that across both benchmarks, the impact of KV-Cloak on inference accuracy is negligible. This confirms our theoretical design: KV-Cloak is a lossless protection scheme that does not degrade the model’s generation quality or its performance on downstream tasks.
@@
V-C4 Evaluation of Computational Overhead
Integrating a new cryptographic primitive into a highly-optimized system like vLLM presents a significant engineering challenge. Therefore, to accurately assess the performance overhead of KV-Cloak and demonstrate its compatibility with PagedAttention, we employ a dual approach: direct latency measurement combined with architectural analysis.
To quantify the computational overhead of KV-Cloak, we measured the inference time for the MMLU benchmark under its protection. We selected this benchmark specifically to evaluate a worst-case scenario: the MMLU task requires applying KV-Cloak’s obfuscation and de-obfuscation to the entire KV-cache of a long input sequence, yet generates only a single output token. This setup maximizes the ratio of cryptographic overhead to inference work. Furthermore, our measurements are conservative as they are based on a serial PyTorch-level implementation without any custom CUDA kernel optimizations, and they do not factor in network communication latency.
TABLE IV: Performance overhead of KV-Cloak (with a block size of 16) on various models during the MMLU benchmark.
Model Inference Time(s)
Plaintext KV-Cloak
LLaMA-7B 10193.1 10649.0 +4.47%
LLaMA-3.2-1B 1956.8 2155.6 +10.16%
LLaMA-3.2-3B-Instruct 4525.7 4821.4 +6.53%
LLaMA-3.1-8B 9797.4 10135.0 +3.45%
LLaMA-3.1-8B-Distilled 9832.8 10297.7 +4.73%
Qwen2.5-Math-7B 9184.7 9452.3 +2.91%
The results of our evaluation across various models are presented in Table IV. The findings show that KV-Cloak introduces a modest average latency overhead of approximately 5%. Notably, we observe a favorable scaling property: the relative overhead decreases as the model’s hidden size increases. In a real-world distributed deployment, the inclusion of network latency between prefill and decode nodes would make the relative impact of KV-Cloak’s computational overhead on the total end-to-end time even lower. These empirical results align with our theoretical analysis, confirming that with standard production optimizations, such as operator fusion, the performance impact of KV-Cloak is expected to be minimal.
Takeaway 7: KV-Cloak introduces negligible computational overhead to the inference pipeline.
VI Conclusion

### v3

@@ V-D Implicit Obfuscation via Operator Fusion
TABLE II: Comparison of defense mechanisms (KV-Cloak vs. DP) against input reconstruction attacks on LLaMA-7B, LLaMA-3.2-1B, and LLaMA-3.1-8B-Distilled.
@@
VI Evaluation of KV-Cloak
VI-A Experimental Settings
We perform a systematic evaluation across three dimensions: Security, Model Accuracy, and Performance Overhead. To benchmark the efficacy of KV-Cloak, we compare it against three distinct baselines: (1) a standard, unprotected system (Plaintext), (2) a defense based on differential privacy with Gaussian noise (DP), and (3) a defense via standard cryptographic encryption (AES).
KV-Cloak Configuration. Our implementation involves matrix multiplications with invertible secret matrices $S$ and $M$, and an additive mask $A$. We configure these parameters to balance security and numerical stability:
  • Block Size $b$: To maintain compatibility with the PagedAttention mechanism (e.g., vLLM), we experimented with standard block sizes of $b\in\{16,32,64\}$.
  • Secret Matrices $S,M$: To minimize precision loss during the matrix inversion required for de-obfuscation, we sample $S$ and $M$ strictly from the orthogonal group.
  • Additive Mask $A$ and Padding: To preserve numerical precision, the additive mask $A$ and padding values are magnitude-constrained. We sample elements of $A$ uniformly from $3\theta_{K},4\theta_{K}]$ and set padding to $1.5\theta_{K}$. Here, $\theta_{K}$ denotes the maximum absolute value observed in the K-cache during calibration on an excerpt from “The Bitter Lesson”. This methodology is symmetrically applied to the V-cache.
DP Baseline Configuration. To establish a strong DP baseline, we conducted an ablation study (detailed in Appendix -D) to optimize the trade-off between privacy and utility. We set the clipping threshold to the 50th percentile of the L2 norm distribution observed across the dataset. Based on this, we apply Gaussian noise calibrated for $(\epsilon=10^{8},\delta=10^{-5})$-DP independently to the K/V caches. We explicitly select $\epsilon=10^{8}$ because stricter privacy budgets (e.g., $\epsilon=10^{7}$) resulted in near-zero inference accuracy (see Table VIII).
VI-B Evaluation of Security
@@
Figure 8: Distance distributions of target tokens $d_{\text{target}}$ (orange) versus incorrect tokens $d_{\text{other}}$ (blue) in the Collision Attack. The experiment targets the last-layer KV-cache of LLaMA-3.2-1B using an excerpt from “The Bitter Lesson”. Subplots display distributions under four conditions: (a) Plaintext, (b) KV-Cloak protection, (c) ($10^{7},10^{-5}$)-DP protection, and (d) ($10^{8},10^{-5}$)-DP protection. Vertical lines indicate the heuristic threshold ($3\sigma_{\text{other}}$, green dotted) and the prior-knowledge enhanced threshold ($r=64$, red dashed).
This section evaluates the capability of KV-Cloak against the three proposed reconstruction attacks. We benchmark against the Plaintext baseline and Differential Privacy mechanisms, applying the attacks to the protected KV-cache.
Defense Efficacy Comparison. As summarized in Table II (evaluation on the remaining models can be found in Appendix -E), the Plaintext baseline is highly vulnerable, yielding high attack success rates on all three attacks. In stark contrast, applying KV-Cloak drastically degrades reconstruction quality. The BERTScore of all attack outputs drops to a level consistent with random chance, and the ROUGE-L score falls to nearly 0. These results are statistically indistinguishable from comparing the original input with a random string, demonstrating that semantic reconstruction is entirely disrupted. This proves that KV-Cloak effectively protects the private information within the KV-cache. For the DP baseline, security is heavily dependent on the privacy budget $\epsilon$. With a weak budget of $\epsilon=10^{8}$, while Inversion and Injection attacks are mitigated, the Collision Attack remains effective, recovering substantial private information.
Statistical Indistinguishability Analysis. To understand the root cause of defense failure or success, we analyze the distance distributions of target tokens ($d_{\text{target}}$) versus other tokens ($d_{\text{other}}$) for the Collision Attack on LLaMA-3.2-1B (Fig. 8). Under DP protection, the two distributions remain statistically distinguishable. Consequently, utilizing an enhanced threshold with prior knowledge yields a per-token success rate of 94.84% for ($10^{7},10^{-5}$)-DP, rising to nearly 100% (equivalent to Plaintext) for ($10^{8},10^{-5}$)-DP. Conversely, KV-Cloak achieves distribution collapse: the distributions of $d_{\text{target}}$ and $d_{\text{other}}$ become completely indistinguishable. This eliminates the statistical separability required for the attack, resulting in a $0\%$ success rate.
Robustness against Enhanced Attacks. To assess the robustness of our defense, We further evaluate by subjecting KV-Cloak to the enhanced Collision Attack (utilizing adversarial prior knowledge) across multiple LLMs. As shown in Table II, KV-Cloak demonstrates consistent resilience. Even against this advanced vector, reconstruction accuracy remains near-zero, and outputs are qualitatively equivalent to random noise. This confirms that KV-Cloak effectively neutralizes both algebraic and statistical attack vectors.
Takeaway 5: KV-Cloak completely thwarts all proposed attacks, reducing the quality of reconstructed text to a level statistically indistinguishable from random noise.
VI-C Inference Accuracy
@@
To rigorously assess model fidelity, we simulate a disaggregated inference service employing a prefill-decode architecture. In this pipeline, the KV-cache generated during the prefill phase is secured using either DP or KV-Cloak, transferred, and subsequently utilized by the decode node for token generation. We employ two standard benchmarks to measure utility: MMLU [16, 15] for massive multitask knowledge and SQuAD [29] for reading comprehension.
We evaluated KV-Cloak across all experimental models (results detailed in Table III). The empirical data corroborates our theoretical design: KV-Cloak exhibits zero degradation in model performance. Unlike Differential Privacy, which forces a trade-off between utility and privacy via noise injection, KV-Cloak leverages reversible linear transformations. This ensures that the de-obfuscated attention states are mathematically identical to the plaintext baseline. Consequently, the impact on inference accuracy is negligible across both benchmarks, confirming that KV-Cloak is a practically lossless solution.
@@
VI-D Evaluation of Performance Overhead
Figure 9: Micro-benchmark of latency overhead (ms/GB) versus KV-cache size on LLaMA-3.1-8B.
Computational Overhead. To accurately assess the upper bound of performance impact, we employ a conservative, worst-case micro-benchmark simulating a prefill-decode split. Measurements are conducted on a serial PyTorch implementation without custom CUDA kernels, isolating the algorithmic overhead. As illustrated in Figure 9, latency scales linearly. Standard AES encryption incurs a prohibitive cost of 3020.9 ms/GB, nearly rivaling the prefill latency (3438.5 ms/GB) itself. In stark contrast, KV-Cloak with operator fusion introduces a negligible overhead of just 15.41 ms/GB. This constitutes merely 0.45% of the prefill latency, ranking second only to the insecure DP scheme (4.88 ms/GB). In practical deployments, unavoidable network latency would further mask this minimal computation cost, confirming the efficiency of our design.
Storage Overhead. The storage Overhead for KV-Cloak’s secret matrices ($S,M,A$) is negligible. The total size, $2n_{\text{kv}}l(b^{2}+\frac{3}{2}d+1)$, is orders of magnitude smaller than the protected cache. Quantitatively, for LLaMA-3.1-8B (using block size $b=16$), the overhead is merely 898 KB; even for a 1T-parameter model like Ling-1T [18] (scaling to $b=64$), it remains only 20.9 MB. This MB-scale footprint allows all cryptographic keys to securely reside within the limited memory of TEEs with negligible management overhead.
Takeaway 7: KV-Cloak introduces negligible computational and storage overhead to the inference pipeline.
VII Discussion and Future Work
While KV-Cloak provides a solid foundation for protecting the KV-cache, we identify several limitations, which in turn open up exciting avenues for future research.
Key Management and Hardware Security Integration. The security of KV-Cloak fundamentally relies on the confidentiality of its (megabyte-scale) key matrices. Currently, we assume these secrets are protected via TEEs. Future work should explore the deep integration of KV-Cloak with hardware-level security mechanisms, such as TEEs or confidential GPUs, to construct a more robust defense-in-depth architecture. Furthermore, to counter long-term cryptanalysis in persistent services, developing efficient, low-latency key rotation protocols is essential.
Performance Optimization via Co-Design. Although our evaluation shows minimal overhead, hyper-scale deployment demands further optimization. At the software level, latency can be masked through the asynchronous generation of One-Time Pad (OTP) matrices, decoupling security operations from the critical inference path. At the hardware level, an algorithm-hardware co-design approach—implementing dedicated GPU intrinsics for block-wise permutation and matrix multiplication—could render the obfuscation cost virtually transparent.
Adaptation to Quantized Models. Our current prototype targets floating-point models. As the industry pivots toward integer quantization (e.g., INT8 or INT4) for efficiency, extending KV-Cloak is a priority. This necessitates designing new, lossless reversible transformations suitable for discrete data types, potentially based on mathematical structures like modular arithmetic.
VIII Conclusion